at the conclusion of a soccer game whose two teams each include 11 players, each player on the winning team "gave five" to (slapped hands) each player on the losing team. each player on the winning team also gave five to each other player on the winning team. how many fives were given?
Answer by edjones(8007)

11^2=121 winners to losers
10+9+8+...1=55 winners to winners.
121+55=176
.
Ed